# Hamiltonian stochastic Lie systems and applications

### v2

@@ 3 Superposition rules and stochastic Lie systems
Let us study the superposition rule notion for stochastic differential equations and the characterisation of stochastic differential equations admitting a superposition rule. This will lead to review and slightly correct some mistakes in the main theorem in [39].
@@
A superposition rule for a Stratonovich stochastic differential equation of the form (2.3) on a manifold $M$ is a function $\Phi:M^{m+1}\to M$ such that, for a generic set $\Gamma_{1},\ldots,\Gamma_{m}:\mathbb{R}_{+}\times\Omega\rightarrow M$, of particular solutions of (2.3), the general solution $\Gamma$ to (2.3) reads
@@
More explicitly, if $Y$ be a vector field on a manifold $M$. The diagonal prolongation of $Y$ to $M^{k}$ is the vector field
@@
The diagonal prolongation of a vector field can be extended to $t$-dependent vector fields on $M$, namely mappings $X\colon\mathbb{R}\times M\rightarrow\mathrm{T}M$ such that $X(t,\cdot)$ is a standard vector field $M$, by assuming that the diagonal extension to $M^{k}$ of the $t$-dependent vector field $X$ on $M$ is the $t$-dependent vector field $\widetilde{X}^{[k]}$ on $M^{k}$ whose value for every fixed valued of $t$, let us say $\widetilde{X}^{[k]}_{t}$, is the diagonal prolongation to $M^{k}$ of the vector field $X_{t}$. The space of diagonal prolongations of vector fields in $\mathfrak{X}(M)$ to $M^{k}$ form a Lie subalgebra of $\mathfrak{X}(M^{k})$. In fact, the mapping $X\in\mathfrak{X}(M)\mapsto{X}^{[k]}\in\mathfrak{X}(M^{k})$ is a Lie algebra morphism, i.e. it is a linear mapping such that
$[Y_{1},Y_{2}]^{[k]}=\left[{Y}^{[k]}_{1},{Y}^{[k]}_{2}\right],\qquad\forall Y_{1},Y_{2}\in\mathfrak{X}(M)\,.$ (3.3)
Let us solve a mistake in the proof of the stochastic Lie theorem in [39]. The issue appears in the direct part of the statement [39, p. 215, Theorem 3.1]. In particular, this makes [39, Remark 3.3.(i)] incorrect. For the sake of completeness, we will state that part of the work adopting our notation and writing in full the contents referenced by the labels used in [39]. We hereafter refer to
“Moreover, the fact
@@
and the hypothesis on $\{Y_{1},\dotsc,Y_{\kappa}\}$ being in involution [^2] It is really meant that the distribution spanned by $Y_{1},\ldots,Y_{\kappa}$ is involutive and $Y_{1},\ldots,Y_{\kappa}$ are linearly independent over $\mathbb{R}$. imply by the classical Frobenius Theorem that
@@
which is false. Let us explain the mistake and build a counterexample. Our counterexample will be relevant as it will show that certain stochastic differential equations in the Itô approach are not stochastic Lie systems and do not accept a superposition rule.
The problem relies on the fact that if $Y_{1},\dotsc,Y_{\kappa}$ span an involutive distribution on $M$, then the diagonal prolongations $\widetilde{Y}^{[k+1]}_{1},\ldots,\widetilde{Y}^{[k+1]}_{\kappa}$ do not span the distribution given by the diagonal prolongations of the distribution spanned by $Y_{1},\ldots,Y_{\kappa}$.
For instance, consider the two vector fields on =$\mathbb{R}_{\circ}=\mathbb{R}\setminus\{0\}$ given by
@@
that span an involutive distribution on =$\mathbb{R}_{\circ}=\{x\in\mathbb{R}:x\neq 0\}$. Notwithstanding, their diagonal prolongations, ${Y}^{[s]}_{1},{Y}^{[s]}_{2}$, to $\mathbb{R}_{\circ}^{s}$ , with $s>2$, do not need to span an involutive distribution. We hereafter consider $s>2$. Indeed, one has the diagonal prolongations
@@
on $\mathbb{R}^{s}_{\circ}$. Their successive commutators become, as stated in [39], diagonal prolongations of an element of the involutive distribution spanned by $Y_{1},Y_{2}$, namely $\mathrm{T}\mathbb{R}_{\circ}$. In particular,
@@
for $k=1,2,\ldots$ Notwithstanding, any $\operatorname{ad}_{{Y}^{[s]}_{1}}^{k}{Y}^{[s]}_{2}$, with $k\in\{2,3,4,\ldots\}$, does not take values in the distribution $\mathcal{D}=\left\langle{Y}^{[s]}_{1},{Y}^{[s]}_{2}\right\rangle$ at a generic point of $\mathbb{R}_{\circ}^{s}$. Even worse, the smallest (in the sense of inclusion) involutive generalised distribution containing ${Y}^{[s]}_{1},{Y}^{[s]}_{2}$ spans the whole tangent space to $\mathbb{R}_{\circ}^{s}$ at almost every point and a superposition rule for a system described by a generic combination of $Y_{1},Y_{2}$ does not exist as it must be constructed from the non-constant first integrals of the vector fields taking values in an involutive distribution containing ${Y}^{[s]}_{1},{Y}^{[s]}_{2}$ for some $s>2$ (see the proof for the Lie theorem in [18, 20, 23, 39]). In fact, the vector fields ${Y}^{[s]}_{\mu}=\sum_{a=1}^{s}x_{(a)}^{\mu+1}\frac{\partial}{\partial x_{(a)}}$, with $\mu=1,\ldots,s$, on $\mathbb{R}_{\circ}^{s}$ are linearly independent almost everywhere. To verify this fact, it is enough to see that the determinant of their coefficients in the basis $\frac{\partial}{\partial x_{(a)}}$, with $a=1,\ldots,s$, read
@@
which makes that the smallest involutive generalised distribution containing ${Y}^{[s]}_{1},{Y}^{[s]}_{2}$ on $\mathbb{R}_{\circ}^{s}$ to be equal to $\mathrm{T}\mathbb{R}_{\circ}^{s}$ almost everywhere. This makes the existence of common non-constant first integrals for ${Y}^{[s]}_{1},{Y}^{[s]}_{2}$, which will be common non-contact first integrals for all $Y^{[s]}_{1},Y^{[s]}_{2},\ldots$ impossible.
The previous counterexample is very important as it concerns the stochastic generalisations of the so-called Abel equations (see [15] and references therein). Hence, the mistake in [39] has potential practical consequences. In particular, note that the SIS model in the Itô form (2.11), for a non-constant function $\sigma(t)$, takes the Stratonovich form (2.12). If one tries to write the $t$-dependent coefficient for $\delta t$ as a linear combination with $t$-dependent functions of a family of vector fields on $\mathbb{R}$ spanning a finite-dimensional Lie algebra, one finds that, for a generic $t$-dependent function $\sigma(t)$, one has to obtain a finite-dimensional Lie algebra of vector fields on $\mathbb{R}$ containing $x^{2}\frac{\partial}{\partial x},x^{3}\frac{\partial}{\partial x}$, which is impossible as shown above.
Despite the above mistake, the stochastic Lie–Scheffers theorem in [39] can be corrected by assuming that the initial family of vector field $Y_{1},\ldots,Y_{r}$ close an $r$-dimensional Lie algebra. Note that the assumption that $Y_{1},\ldots,Y_{r}$ are linearly independent over $\mathbb{R}$ is necessary. Otherwise, e.g. if $Y_{r-1}=Y_{r}$, one obtains that the vector fields $Y_{1}^{[m]},\ldots,Y_{r}^{[m]}$ are always linearly dependent and Lemma 3.2 of [39] cannot be applied in [39, p. 918]. On the other hand, if $Y_{1},\ldots,Y_{r}$ are linearly independent, their diagonal prolongations become linearly independent at a generic point exactly when $m$ is such that they span a distribution of rank $r$ at a generic point (cf. [20]). Then, Lemma 3.2 can be applied normally as accomplished in [39, p. 918].
@@
Let us prove the part of the direct implication of the theorem in [39, Theorem 3.1] that requires some comments due to our correction. One has that there always exists a number $m$ such that the diagonal prolongation of a basis $Y_{1},\ldots,Y_{r}$ of the Vessiot–Guldberg Lie algebra of the Lie system reaches rank $r$ at a generic point. It was proved in [20] that this happens exactly when the diagonal prolongations span a distribution of rank $r$ almost everywhere. Let be $\mathcal{D}_{0}$ the distribution spanned by vector fields $Y^{[m+1]}_{1},\ldots,Y^{[m+1]}_{r}$. Then $\mathcal{D}_{0}$ has constant rank on an open subset of $M^{m+1}$. Since the Lie brackets $[Y_{i},Y_{j}]$ are linear combinations of the vector fields of $V$, with constant coefficients, then their diagonal prolongations to $M^{m+1}$ are linear combinations, with constant coefficients, of the $Y^{[m+1]}_{1},\ldots,Y^{[m+1]}_{r}$ and $\mathcal{D}_{0}$ is integrable. Then, one can extend $\mathcal{D}_{0}$ to an integrable distribution $\mathcal{D}$ of corank $n$ in $M^{m+1}$, i.e. $\mathcal{D}_{0}\subset\mathcal{D}$ on every point of an open neighbourhood of a point in $M^{m+1}$. This can always be done in such a manner that the leaves of $\mathcal{D}$ project diffeomorphically onto an open subset of $M^{m}$ via the canonical projection $\pi:(\Gamma_{(1)},\ldots,\Gamma_{(m+1)})\in M^{m+1}\mapsto(\Gamma_{(1)},\ldots,\Gamma_{(m)})\in M^{m}$. The obtained expression is a local superposition rule for (3.4) as shown in [39, Theorem 3.1] in basis of [39, Proposition 2.4]. ∎
@@ 4 Hamiltonian stochastic Lie systems
Since $[X_{1},X_{2}]=-X_{2},$ the vector fields $X_{1},X_{2}$ span a non-Abelian two-dimensional Lie algebra $\mathfrak{h}_{2}$. Hence, system (4.1) is a stochastic Lie system.
@@
where we consider solutions $\Gamma:(t,\omega_{p})\in\mathbb{R}_{+}\times\Omega\mapsto(S,I)\in\mathcal{U}=\{(S,I)\in\mathbb{R}^{2}:SI\neq 0\}$. In the particular case of $\Lambda=0$, one can consider the locally defined symplectic form, away from $(0,0)$, of the form
@@
where $\xi_{1},\xi_{2}$ are local coordinates such that $\bar{X}_{1}=\frac{\partial}{\partial\xi_{1}}$ and $\bar{X}_{2}=\frac{\partial}{\partial\xi_{2}}$. The existence of $\xi_{1},\xi_{2}$ is due to the fact that $\bar{X}_{1}\wedge\bar{X}_{2}$ does not vanish and $[\bar{X}_{1},\bar{X}_{2}]=0$ away from $(0,0)$. Note that system (4.2) can be generalised to consider $t$-dependent coefficients. the stochastic differential equation on $\mathbb{R}^{n}$ with a Wiener process $W_{1}$ (see [35, 36, 57] for analogues on $\mathbb{R}$) in Stratonovich form given by
$\delta\begin{pmatrix}x_{i}\\ y_{i}\end{pmatrix}=\begin{pmatrix}\lambda&a\\ b&-\lambda\end{pmatrix}\begin{pmatrix}x_{i}\\ y_{i}\end{pmatrix}\delta t+\begin{pmatrix}\lambda^{\prime}&a^{\prime}\\ b^{\prime}&-\lambda^{\prime}\end{pmatrix}\begin{pmatrix}x_{i}\\ y_{i}\end{pmatrix}\circ\delta W_{1}\,,\qquad i=1,\ldots,n\,,$ (4.4)
which is defined on $\mathrm{T}^{\ast}\mathbb{R}^{n}$. This recovers Example 2.3, namely Eq. (2.9), by taking $k=0$ and $\sigma=0$ as a special case. It can be proved that (2.9) is not Hamiltonian relative to any symplectic form on $\mathbb{R}^{2}$ for arbitrary $t$-dependent coefficients as it leads to a Vessiot–Guldberg Lie algebra on the plane related to $\mathfrak{gl}_{2}$ [23]. But (4.4) can now be related to a Vessiot–Guldberg Lie algebra $V$ on $\mathrm{T}^{*}\mathbb{R}^{n}$ spanned by
$X_{1}=\sum_{i=1}^{n}y_{i}\frac{\partial}{\partial x_{i}}\,,\qquad X_{2}=\frac{1}{2}\sum_{i=1}^{n}\left(x_{i}\frac{\partial}{\partial x_{i}}-y_{i}\frac{\partial}{\partial y_{i}}\right)\,,\qquad X_{3}=-\sum_{i=1}^{n}x_{i}\frac{\partial}{\partial y_{i}}\,.$
@@
$[X_{1},X_{2}]=X_{1}\,,\qquad[X_{1},X_{3}]=X_{2}\,,\qquad[X_{2},X_{3}]=X_{3}\,,$
this Lie algebra is isomorphic to $\mathfrak{sl}_{2}$ and it is the diagonal prolongation to $(\mathrm{T}^{*}\mathbb{R})^{n}$ of a Vessiot–Guldberg Lie algebra isomorphic to $\mathfrak{sl}_{2}$ on $\mathrm{T}^{*}\mathbb{R}$ (see [6, 23]). Moreover, it is known to be a Lie algebra of Hamiltonian vector fields relative to the symplectic form
@@
$h_{1}=\sum_{i=1}^{n}\frac{y_{i}^{2}}{2}\,,\qquad h_{2}=\frac{1}{2}\sum_{i=1}^{n}x_{i}y_{i}\,,\qquad h_{3}=\sum_{i=1}^{n}\frac{x_{i}^{2}}{2}\,.$
@@
$\mathfrak{S}(x,y,t,W)=\left(\lambda X_{2}+aX_{1}-bX_{3},\lambda^{\prime}X_{2}+a^{\prime}X_{1}-b^{\prime}X_{3}\right),$
@@
respectively. It is standard to call $\widetilde{h}_{1}$ the Hamiltonian of the stochastic differential equation.
Figure 1: These are two representations of the evolution in terms of the time of a particular solution to the Hamiltonian Lie system $\delta y=-x\delta t-Ax\delta W,\delta x=y\delta t+Ay\delta W$, with initial condition (.$(.2,0)$ a semi-martingale $W$, and a parameter $A$. The symplectic structure is $\omega=\mathrm{d}x\wedge\mathrm{d}y$. The system has a strong constant of motion $x^{2}+y^{2}$. The constant of the motion is always conserved, but solutions jump back and forward relative to the deterministic solution.

### v3

@@ 3 Superposition rules and stochastic Lie systems
Let us study the superposition rule notion for stochastic differential equations and the characterisation of stochastic differential equations admitting a superposition rule. This will lead us to review and slightly correct some mistakes in the main theorem in [48].
@@
A superposition rule for a Stratonovich stochastic differential equation of the form (2.3) on a manifold $M$ is a function $\Phi:M^{m+1}\to M$ such that, for a generic set $\Gamma_{1},\ldots,\Gamma_{m}:\mathbb{R}_{+}\times\Omega\rightarrow M$, of particular solutions of (2.3), the general solution $\Gamma$ to (2.3) takes the form
@@
More explicitly, let $Y$ be a vector field on a manifold $M$. The diagonal prolongation of $Y$ to $M^{k}$ is the vector field
@@
The diagonal prolongation of a vector field can be extended to $t$-dependent vector fields on $M$, namely mappings $X\colon\mathbb{R}\times M\rightarrow\mathrm{T}M$ such that $X(t,\cdot)$ is a standard vector field on $M$, by assuming that the diagonal extension to $M^{k}$ of the $t$-dependent vector field $X$ on $M$ is the $t$-dependent vector field $\widetilde{X}^{[k]}$ on $M^{k}$ whose value for every fixed valued of $t$, let us say $\widetilde{X}^{[k]}_{t}$, is the diagonal prolongation to $M^{k}$ of the vector field $X_{t}$. The space of diagonal prolongations of vector fields in $\mathfrak{X}(M)$ to $M^{k}$ form a Lie subalgebra of $\mathfrak{X}(M^{k})$. In fact, the mapping $X\in\mathfrak{X}(M)\mapsto{X}^{[k]}\in\mathfrak{X}(M^{k})$ is a Lie algebra morphism, i.e. it is a linear mapping such that
$Y_{1},Y_{2}]^{[k]}=\left[{Y}^{[k]}_{1},{Y}^{[k]}_{2}\right],\qquad\forall Y_{1},Y_{2}\in\mathfrak{X}(M)\,.$ (3.3)
Let us solve a mistake in the proof of the stochastic Lie theorem in [48]. The issue appears in the direct part of the statement [48, p. 215, Theorem 3.1]. In particular, this makes [48, Remark 3.3.(i)] incorrect. To start with, just recall a couple of notions on differential geometry to make our presentation more accessible for researchers working on stochastic differential equations or applications.
A generalised distribution $D$ on a manifold $M$ is a correspondence mapping every point $p\in M$ to a subspace $D_{p}\subset T_{p}M$. A generalised distribution $D$ is smooth if, around every point $p\in M$, there exists an open subset $U\ni p$ and vector fields $X_{1},\ldots,X_{s}$ on $U$ such that $D_{p^{\prime}}=\langle X_{1}(p^{\prime}),\ldots,X_{s}(p^{\prime})\rangle$ for every $p^{\prime}\in U$. It is worth noting that the number $s$ may depend on the point $p$. A generalised distribution $D$ is involutive if for every pair of vector fields $X_{1},X_{2}$ on an open subset $U\subset M$ such that $X_{1}(p),X_{2}(p)\in D_{p}$ for every $p\in U$, we have that $X_{1},X_{2}](p)\in D_{p}$ at every $p\in U$. In other words, a generalised distribution is involutive if the Lie bracket of vector fields taking values in the distribution takes values in the distribution too. More practically, it can be proved that a generalised distribution $D$ on $M$ is involutive if for every point $p\in M$ there exists a family of vector fields $X_{1},\ldots,X_{s}$ spanning the distribution on an open neighbourhood $U$ of $p$, namely $D_{p^{\prime}}=\langle X_{1}(p^{\prime}),\ldots,X_{s}(p^{\prime})\rangle$ for $p^{\prime}\in U$, satisfy that the Lie brackets $X_{i},X_{j}]$, with $1\leq i<j\leq s$, also take values in $D$ on $U$, namely $X_{i},X_{j}](p^{\prime})\in D_{p^{\prime}}$ for every $p^{\prime}\in U$ and $1\leq i<j\leq s$. In this work, we have said ‘distributions’ instead of ‘generalised distributions’ to simplify our terminology, as commonly done in the literature. Moreover, all inspected distributions are smooth.
For the sake of completeness, we will state that part of the work adopting our notation and writing in full the contents referenced by the labels used in [48, Theorem 3.1]. We hereafter refer to
“Moreover, the fact that one has the general property
@@
and the hypothesis on $\{Y_{1},\dotsc,Y_{\kappa}\}$ that they span an involutive distribution [^1] In [48, Theorem 3.1], it is stated that $\{Y_{1},\ldots,Y_{\kappa}\}$ span an involutive distribution, i.e. $Y_{i},Y_{j}]=\sum_{k=1}^{\kappa}f_{ijk}Y_{k}$ for suitable functions $f_{ijk}$ with $i,j,k=1,\ldots,r$. The proof also notes that $\{Y_{1},\ldots,Y_{r}\}$ are ‘in involution,’ again meaning they span an involutive distribution. This terminology may be misleading, as in some references ‘vector fields in involution’ is used to mean that the vector fields commute. imply by the classical Fröbenius theorem that
@@
which is incorrect. Let us explain the mistake and provide a counterexample. Our counterexample will be relevant because it will show that certain stochastic differential equations in the Itô approach are not stochastic Lie systems and do not admit a superposition rule.
The problem relies on the fact that if $Y_{1},\dotsc,Y_{\kappa}$ span an involutive distribution on $M$, then the diagonal prolongations $\widetilde{Y}^{[k+1]}_{1},\ldots,\widetilde{Y}^{[k+1]}_{\kappa}$ do not need to span the distribution given by the diagonal prolongations of the vector fields taking values in the distribution spanned by $Y_{1},\ldots,Y_{\kappa}$. In fact, as shown next, the distribution spanned by $\widetilde{Y}^{[k+1]}_{1},\ldots,\widetilde{Y}^{[k+1]}_{\kappa}$ need not be involutive at all.
For instance, consider the two vector fields on $\mathbb{R}_{\circ}=\mathbb{R}\setminus\{0\}$ given by
@@
that span an involutive distribution on $\mathbb{R}_{\circ}=\{x\in\mathbb{R}|x\neq 0\}$. Indeed, both vector fields span a distribution $\mathcal{D}_{x}=\langle x^{2}\partial/\partial x,x^{3}\partial/\partial\rangle=\mathrm{T}_{x}\mathbb{R}_{\circ}$ for $x\in\mathbb{R}_{\circ}$ and every two vector fields taking values in $\mathrm{T}\mathbb{R}_{\circ}$ have a Lie bracket contained in $\mathrm{T}\mathbb{R}_{\circ}$. Notwithstanding, their diagonal prolongations, ${Y}^{[s]}_{1},{Y}^{[s]}_{2}$, to $\mathbb{R}_{\circ}^{s}$ with $s>2$ do not need to span an involutive distribution. Indeed, one has the diagonal prolongations
@@
on $(\mathbb{R}_{\circ})^{s}$. Their successive commutators become, as stated in [48], diagonal prolongations of an element of the involutive distribution spanned by $Y_{1},Y_{2}$, namely $\mathrm{T}\mathbb{R}_{\circ}$. In particular,
@@
for $k=1,2,\ldots$ Notwithstanding, $\operatorname{ad}_{{Y}^{[s]}_{1}}^{k}{Y}^{[s]}_{2}$, with $k\in\{2,3,4,\ldots\}$, does not take values in the distribution $\mathcal{D}=\left\langle{Y}^{[s]}_{1},{Y}^{[s]}_{2}\right\rangle$ at a generic point of $\mathbb{R}_{\circ}^{s}$. Even worse, the smallest (in the sense of inclusion) involutive distribution containing ${Y}^{[s]}_{1},{Y}^{[s]}_{2}$ spans the whole tangent space to $\mathbb{R}_{\circ}^{s}$ at almost every point and a superposition rule for a system described by a generic combination of $Y_{1},Y_{2}$ does not exist as it must be constructed from the non-constant first integrals of the vector fields taking values in an involutive distribution containing ${Y}^{[s]}_{1},{Y}^{[s]}_{2}$ for some $s>2$ (see the proof for the Lie theorem in [19, 23, 26, 48]). In fact, the vector fields ${Y}^{[s]}_{\mu}=\sum_{a=1}^{s}x_{(a)}^{\mu+1}\frac{\partial}{\partial x_{(a)}}$, with $\mu=1,\ldots,s$, on $\mathbb{R}_{\circ}^{s}$ are linearly independent almost everywhere. To verify this fact, it is enough to see that the determinant of their coefficients in the basis $\frac{\partial}{\partial x_{(a)}}$, with $a=1,\ldots,s$, read
@@
which causes the smallest involutive generalised distribution containing ${Y}^{[s]}_{1},{Y}^{[s]}_{2}$ on $\mathbb{R}_{\circ}^{s}$ to be equal to $\mathrm{T}\mathbb{R}_{\circ}^{s}$ almost everywhere. This makes the existence of common non-constant first integrals for ${Y}^{[s]}_{1},{Y}^{[s]}_{2}$, which will be common non-constant first integrals for all $Y^{[s]}_{1},Y^{[s]}_{2}$$\ldots$ impossible.
The previous counterexample is very important as it concerns the stochastic generalisations of the so-called Abel equations (see [23] and references therein). Hence, the mistake in [48] has potential practical consequences. Moreover, recall that the SIS model in the Itô form (2.12), for a non-constant function $\sigma(t)$, takes the Stratonovich form (2.13). We already showed that if one tries to write the $t$-dependent coefficient of the related Stratonovich operator for $\delta t$ as a linear combination with $t$-dependent functions of a family of vector fields on $\mathbb{R}$ spanning a finite-dimensional Lie algebra, one finds that, for a generic $t$-dependent function $\sigma(t)$, one has to obtain a finite-dimensional Lie algebra of vector fields on $\mathbb{R}$ containing $x^{2}\frac{\partial}{\partial x},x^{3}\frac{\partial}{\partial x}$, which is impossible as shown in Example 2.1.
Despite the above mistake, the stochastic Lie–Scheffers theorem in [48] can be corrected by assuming that the initial family of vector field $Y_{1},\ldots,Y_{r}$ close an $r$-dimensional Lie algebra. Note that the assumption that $Y_{1},\ldots,Y_{r}$ are linearly independent over $\mathbb{R}$ is necessary. Otherwise, e.g. if $Y_{r-1}=Y_{r}$, one obtains that the vector fields $Y_{1}^{[m]},\ldots,Y_{r}^{[m]}$ are always linearly dependent and Lemma 3.2 of [48] cannot be applied in [48, p. 918]. On the other hand, if $Y_{1},\ldots,Y_{r}$ are linearly independent, their diagonal prolongations become linearly independent at a generic point exactly when $m$ is such that they span a distribution of rank $r$ at a generic point (cf. [23]). Then, Lemma 3.2 can be applied normally as in [48, p. 918].
@@
Let us prove the part of the direct implication of [48, Theorem 3.1] that requires some comments in light of our correction. There always exists a number $m$ such that the diagonal prolongation of a basis $Y_{1},\ldots,Y_{r}$ of the Vessiot–Guldberg Lie algebra of the Lie system to $M^{m}$ reaches rank $r$ at a generic point. It was proved in [23] that this happens exactly when the diagonal prolongations span a distribution of rank $r$ almost everywhere. Let be $\mathcal{D}_{0}$ the distribution spanned by vector fields $Y^{[m+1]}_{1},\ldots,Y^{[m+1]}_{r}$. Then $\mathcal{D}_{0}$ has constant rank on an open subset of $M^{m+1}$. Since the Lie brackets $Y_{i},Y_{j}]$ are linear combinations of the vector fields of $V$, with constant coefficients, then their diagonal prolongations to $M^{m+1}$ are linear combinations, with constant coefficients, of the $Y^{[m+1]}_{1},\ldots,Y^{[m+1]}_{r}$ and $\mathcal{D}_{0}$ is integrable. Then, one can extend $\mathcal{D}_{0}$ to an integrable distribution $\mathcal{D}$ of corank $n$ in $M^{m+1}$, i.e. $\mathcal{D}_{0}\subset\mathcal{D}$ on every point of an open neighbourhood of a point in $M^{m+1}$. This can always be done in such a manner that the leaves of $\mathcal{D}$ project diffeomorphically onto an open subset of $M^{m}$ via the canonical projection $\pi:(\Gamma_{(1)},\ldots,\Gamma_{(m+1)})\in M^{m+1}\mapsto(\Gamma_{(1)},\ldots,\Gamma_{(m)})\in M^{m}$. The obtained expression is a local superposition rule for (3.4) as shown in [48, Theorem 3.1] in basis of [48, Proposition 2.4]. ∎
It is important to highlight that the proof of the stochastic Lie theorem shows, like its non-stochastic analogue, that if the Vessiot–Guldberg Lie algebra has dimension $r$, then $m$, which standard for the number of particular solutions of the associated superposition rule, must fulfils the condition $m\dim M\geq r$. This requirement is crucial to guarantee that the diagonal prolongations $Y^{[m]}_{1},\ldots,Y^{[m]}_{r}$ are linearly independent at a generic point of $M^{m}$. Consequently, a larger Vessiot–Guldberg Lie algebra on a manifold $M$ is associated with a greater number of particular solutions for the superposition rule of the related stochastic Lie system, namely $m$,. On the other hand, larger Vessiot–Guldberg Lie algebras lead to superposition rules for larger families of stochastic Lie systems admitting a common superposition rule (see [23] for an exploration of these facts in the classic non-stochastic context, which is completely analogous).
@@ 4 Hamiltonian stochastic Lie systems
Since $X_{1},X_{2}]=-X_{2},$ the vector fields $X_{1},X_{2}$ span a non-Abelian two-dimensional Lie algebra $\mathfrak{h}_{2}$. Hence, system (4.1) is a stochastic Lie system.
@@
where we consider solutions $\Gamma:(t,\omega_{p})\in\mathbb{R}_{+}\times\Omega\mapsto(S,I)\in\mathcal{U}=\{(S,I)\in\mathbb{R}^{2}:SI\neq 0\}$. In the particular case of $\Lambda=0$, one can consider the locally defined symplectic form, away from points with $(S+I)I\neq 0$, of the form
@@
where $\xi_{1},\xi_{2}$ are local coordinates such that $\bar{X}_{1}=\frac{\partial}{\partial\xi_{1}}$ and $\bar{X}_{2}=\frac{\partial}{\partial\xi_{2}}$. The existence of $\xi_{1},\xi_{2}$ is due to the fact that $\bar{X}_{1}\wedge\bar{X}_{2}$ does not vanish and $\bar{X}_{1},\bar{X}_{2}]=0$ away from points with $(S+I)I\neq 0$. Note that system (4.2) can be generalised to consider $t$-dependent coefficients.
Let us consider the stochastic differential equation on $\mathbb{R}^{n}$ with a Wiener process $W_{1}$ (see [41, 42, 70] for analogues on $\mathbb{R}$) in Stratonovich form given by
$\delta\begin{pmatrix}x_{i}\\ y_{i}\end{pmatrix}=\begin{pmatrix}\lambda(t)&a(t)\\ b(t)&-\lambda(t)\end{pmatrix}\begin{pmatrix}x_{i}\\ y_{i}\end{pmatrix}\delta t+\begin{pmatrix}\lambda^{\prime}(t)&a^{\prime}(t)\\ b^{\prime}(t)&-\lambda^{\prime}(t)\end{pmatrix}\begin{pmatrix}x_{i}\\ y_{i}\end{pmatrix}\circ\delta W_{1}\,,\qquad i=1,\ldots,n\,$, (4.4)
which is defined on $\mathrm{T}^{\ast}\mathbb{R}^{n}$. This recovers subcases of Example 2.4, namely equation (2.9) with $k(t)=\sigma(t)=0$. It can be proved that (2.9) is not Hamiltonian relative to any symplectic form on $\mathbb{R}^{2}$ for arbitrary $t$-dependent coefficients as it leads to a Vessiot–Guldberg Lie algebra on the plane related to $\mathfrak{gl}_{2}$ [26]. But (4.4) can now be related to a Vessiot–Guldberg Lie algebra $V$ on $\mathrm{T}^{*}\mathbb{R}^{n}$ spanned by
$X_{1}=\sum_{i=1}^{n}y_{i}\frac{\partial}{\partial x_{i}}\,,\qquad X_{2}=\frac{1}{2}\sum_{i=1}^{n}\left(x_{i}\frac{\partial}{\partial x_{i}}-y_{i}\frac{\partial}{\partial y_{i}}\right)\,,\qquad X_{3}=-\sum_{i=1}^{n}x_{i}\frac{\partial}{\partial y_{i}}\,.$ (4.5)
@@
$X_{1},X_{2}]=X_{1}\,,\qquad[X_{1},X_{3}]=X_{2}\,,\qquad[X_{2},X_{3}]=X_{3}\,$,
this Lie algebra is isomorphic to $\mathfrak{sl}_{2}$ and it is the diagonal prolongation to $(\mathrm{T}^{*}\mathbb{R})^{n}$ of a Vessiot–Guldberg Lie algebra isomorphic to $\mathfrak{sl}_{2}$ on $\mathrm{T}^{*}\mathbb{R}$ (see [26]). Moreover, it is known to be a Lie algebra of Hamiltonian vector fields relative to the symplectic form
@@
$h_{1}=\sum_{i=1}^{n}\frac{y_{i}^{2}}{2}\,,\qquad h_{2}=\frac{1}{2}\sum_{i=1}^{n}x_{i}y_{i}\,,\qquad h_{3}=\sum_{i=1}^{n}\frac{x_{i}^{2}}{2}\,.$ (4.6)
@@
$\mathfrak{S}(x,y,t,W_{1})=\left(\lambda(t)X_{2}+a(t)X_{1}-b(t)X_{3},\lambda^{\prime}(t)X_{2}+a^{\prime}(t)X_{1}-b^{\prime}(t)X_{3}\right),$
@@
respectively. It is standard to call $\widetilde{h}_{1}$ the Hamiltonian of the stochastic differential equation. We call $\langle h_{1},\ldots,h_{r}\rangle$ a Lie–Hamilton Lie algebra of a Hamiltonian stochastic Lie system.
Figure 1: These are two representations of the evolution in terms of the time of a particular solution to the Hamiltonian Lie system $\delta y=-x\delta t-Ax\delta W,\delta x=y\delta t+Ay\delta W$, with initial condition $(.2,0)$ a semi-martingale $W$, and a parameter AA. The symplectic structure is $\omega=\mathrm{d}x\wedge\mathrm{d}y$. The system has a strong constant of motion $x^{2}+y^{2}$. The constant of motion is always conserved, but solutions jump back and forward relative to the deterministic solution.